OpenStudy (anonymous):

3f(x-2)+f(2-x)=5x f(x)=?

OpenStudy (zzr0ck3r):

what class is this for?

OpenStudy (zzr0ck3r):

Sorry I have no idea. @Callisto please tag me if you see this get solved.

ganeshie8 (ganeshie8):

\(x = 2\) gives \(3f(0) + f(0) = 10\implies f(0) = 5/2\) differentiating we get \(f'(0) = 5/2\) and \(f^n(0) = 0\) for \(n\gt 2\) so \(f(x) \) is a linear polynomial that passes through \((0, 5/2)\) with a slope of \(5/2\) : \[f(x) = \frac{5}{2}x + \frac{5}{2}\]

OpenStudy (zzr0ck3r):

why is f'(0) = 5/2?

ganeshie8 (ganeshie8):

3f(x-2)+f(2-x)=5x differentiateing 3f'(x-2) - f'(2-x) = 5 plugin x=2

OpenStudy (zzr0ck3r):

right

OpenStudy (zzr0ck3r):

Very cool! nice work

ganeshie8 (ganeshie8):

works assuming f(x) is differentiable etc...

OpenStudy (zzr0ck3r):

yeah, very clever